ADNAN MUHAMMED

Grade 12,

A stone is projected horizontally with 20m/s from top of a building . Calculate its position and velocity after 3 seconds neglecting air resistance

A stone is projected horizontally with 20m/s from top of a building . Calculate its position and velocity after 3 seconds neglecting air resistance
 
 

Grade:11

1 Answers

Arun
5 years ago
Since, body is projected horizontally with a velocity of 20m/s.
ux = 20m/s
ax = 0
So, vx = 20m/s for all time.
Along vertical axis,
uy = 0
t = 5
ay = 10m/s2
vy =?
vy = uy +ayt
=> vy = 0 + 10*5= 50m/s
So, magnitude of velocity of body after 5s
√{(50)²+(20)²}m/s
=√2900 m/s
=53.851 m/s
= 54 m/s (approx)